مفصل 3.18 Calculate the unknown temperature in each of the following: , 25°C = | b. 155 °C = °F d. 224K = °C e 15°C = _K
Answer #1

Do rate if you understand

Using the Formola F-32 - 32 ) OF = 45+32= 77 of 31 5 55 = 1-32 ºf 3 32+ 279 = 311] 0 °c = 5(-25-32] = $31.67°) d 244 K = 244